# ChatThero: An LLM-Supported Chatbot for Behavior Change and Therapeutic Support in Addiction Recovery

### v1

@@ 4.1 Dataset
We constructed a large-scale synthetic dataset to simulate realistic therapeutic dialogues related to substance use and recovery. The initial data comprises anonymized and carefully filtered Reddit posts from addiction recovery communities (r/leaves, r/addiction), involving 57,471 unique authors, averaging 18.25 posts per author, with 2.13 main posts explicitly addressing substance use challenges. To protect privacy and adhere to ethical standards, all personally identifiable information (PII) was rigorously removed, verified via manual checks.
@@ 4.2 Model Baseline
4.3 Clinical Scenarios Design
In our experimental setting, we categorize virtual patients into three difficulty levels: Easy, Medium, and Hard. These levels correspond to the anticipated difficulty of persuasion success and the severity of the patient’s substance use dependence.
  • Easy: Patients are generally receptive to therapeutic interventions and demonstrate willingness to follow treatment plans.
  • Medium: Patients exhibit partial resistance, potentially rejecting some proposed strategies, requiring the therapist agent to dynamically adapt persuasion approaches.
  • Hard: Patients show deeply entrenched pessimism about treatment, doubting both their ability to recover and the effectiveness of therapeutic interventions.
We simulate conversations with 300 virtual patients in total, consisting of 100 patients per difficulty level. Each patient is initialized with a unique clinical profile describing their psychological characteristics, substance use history, and motivational barriers.
Clinicians were asked to role-play as patients and engage in full conversations with the model. These interactions typically lasted around 45 minutes and comprised approximately 60 utterances. Based on these observations, we set the maximum conversation length to 60 utterances to reflect a realistic upper bound for therapist-patient interactions. Dialogues are designed to terminate early if the patient expresses resolution of their concerns or demonstrates sufficient confidence in overcoming addiction. Thus, reaching the 60-utterance limit indicates that the patient remains unconvinced or unresolved by the end of the session. This setup ensures comparability across patients. During evaluation, we analyze the diversity and deployment of persuasive strategies employed by the therapy agent. The strategy pool is shown in the Table 7. We focus on quantifying both the frequency and variety of strategies used within each conversation, and compare strategy usage patterns across patient difficulty levels.
4.4 Human evaluation.
To systematically assess the quality of doctor-patient conversations generated by different models, we designed a structured evaluation protocol using a five-dimensional scoring framework  [^2] The expert annotators were two licensed physicians. Both consents were obtained before they began the study, and they were given a \$ 50 gift card for their participation as compensation.. Each human expert evaluator is instructed to act as a professional clinical conversation assessor and assign a score from 1 to 5 (with decimals allowed for finer granularity) on the following dimensions:
  • Responsiveness (R), which captures how effectively the doctor addresses the patient’s concerns, emotions, and questions;
  • Empathy (E), reflecting the degree of emotional sensitivity and support conveyed by the doctor;
  • Persuasive Strategy Appropriateness (P), assessing whether the chosen strategies (e.g., evidence-based reasoning, analogies, motivational interviewing) align with the patient’s level of resistance;
  • Clinical Relevance (C), measuring the accuracy and therapeutic validity of the content within a substance use context; and
  • Behavioral Realism (B), which evaluates whether the overall interaction style, tone, and pacing resemble real-world clinician behavior.
In addition, we also involved experts in the evaluation process. Unlike previous evaluations, the experts conducted comparative assessments—comparing responses from our model against those from other models to calculate the win rate.
4.5 LLM-as-a-Juage
Large language models are increasingly used as automated judges to evaluate text generation, showing strong agreement with human annotators across diverse tasks such as summarization, dialogue, and clinical scenarios Arora et al. (2025); Li et al. (2024a); Achiam et al. (2023); Tu et al. (2025). Following this trend, we implement a GPT-4o-based LLM-as-a-Judge using the same evaluation dimensions and settings as our human assessment. Details and prompts are provided in Table 7 and 12.
@@ 5 Results and Discussion
We present comprehensive results from quantitative and qualitative evaluations, addressing both the robustness of our experimental design and the implications for scalable addiction treatment interventions.
5.1 Persuasion Outcomes Across Difficulty Levels
Model Easy Medium Hard
ChatThero 5.0 / 4.1 4.2 / 3.4 4.1 / 3.2
GPT-4o 5.0 / 3.9 2.8 / 1.8 1.6 / 1.1
GPT-4o-mini 4.9 / 3.8 2.4 / 1.4 1.6 / 1.0
Table 1: Average motivation/confidence scores across models and patient difficulty levels.
Table 1 illustrates the average motivation and confidence scores achieved by various models across patient difficulty levels. While baseline models performed comparably in easy scenarios, substantial gaps emerged with increasing patient resistance. Notably, ChatThero significantly improved outcomes in medium (motivation +1.4, confidence +1.6 vs. GPT-4o) and hard scenarios (motivation +2.5, confidence +2.1 vs. GPT-4o), demonstrating its superior adaptive capabilities in handling challenging patient interactions. These results highlight ChatThero’s strong potential for practical implementation in realistic therapeutic settings.
Figure 3: Motivation, Confidence, and the Impact of Strategic Diversity Across Patient Difficulty Levels and Models. Left panel: Motivation and confidence score trajectories over conversation turns for Easy, Medium, and Hard virtual patients, evaluated across all models. The x-axis denotes the number of dialogue turns; the y-axis shows the average motivation (top) or confidence (bottom) score at the final turn of dialogues ending at each point. Each curve corresponds to a different model, with bubble sizes indicating the frequency of dialogue termination at that turn count. Effective models (notably ChatThero) enable more patients to achieve higher motivation and confidence earlier, leading to earlier and more frequent session resolution—especially for harder patients. Right panel: Relationship between cumulative unique persuasive strategies employed and resulting motivation (top) or confidence (bottom) scores for each model and difficulty group. Solid, dashed, and dotted lines represent Easy, Medium, and Hard patients, respectively. Colors denote different models. ChatThero consistently demonstrates a strong positive correlation between the diversity of strategies used and higher motivation/confidence scores, particularly in Medium and Hard patient scenarios, underscoring the importance of adaptive and varied intervention techniques for effective therapeutic support.
5.2 Trajectory of Persuasion and Early Termination Efficiency
5.3 Quality of Persuasive Communication
Therapy Metric R E P C B W H
GPT-4o 4.68 4.87 4.39 4.47 4.50 65.5% 62%
GPT-4o mini 4.66 4.86 4.38 4.49 4.46 69.4% 71%
Qwen2.5-7B 4.33 4.58 4.02 4.53 4.24 85.3% 90%
Qwen2.5-14B 4.52 4.58 4.28 4.56 4.38 83.3% 89%
Qwen2.5-32B 4.53 4.56 4.25 4.60 4.36 82.4% 85%
LLaMA3.1-8B 4.43 4.62 4.13 4.58 4.34 82.0% 91%
ChatThero-FT 4.81 4.90 4.66 4.58 4.65 - -
ChatThero-DPO 4.85 4.93 4.75 4.61 4.69 42.3% 41%
Table 2: Model performance on five core persuasion dimensions. Win Rate reflects pairwise preference against ChatThero-FT using GPT-4o. Human denotes the proportion of conversations evaluated by human annotators.
Evaluations across five clinical dimensions (Table 2) indicate ChatThero consistently outperformed baseline models, especially regarding empathy (4.93), persuasive strategy appropriateness (4.75), and behavioral realism (4.69). Pairwise preference comparisons further validated the incremental benefit of DPO optimization over supervised fine-tuning (ChatThero-DPO winning 42.3% of comparisons against its SFT counterpart). Human expert ratings strongly correlated with GPT-4o automated evaluations, confirming reliability and reproducibility of our evaluation protocols.
5.4 Qualitative Analysis
Detailed dialogue comparisons between ChatThero and baselines (Tables 1415-18) vividly illustrate ChatThero’s superior adaptive responsiveness, personalized strategic layering, and emotional sensitivity. In contrast, baseline models often demonstrated rigid or generic approaches, insufficiently addressing patient resistance or emotional nuances. Such qualitative insights deepen our understanding of the specific conversational qualities necessary for effective therapeutic interventions.
5.5 Ablation Study
We further analyzed the incremental contributions of our training approach through an ablation study comparing base Qwen-7B, ChatThero-FT, and ChatThero-DPO variants. Results clearly indicate the necessity of DPO-based optimization, which significantly enhances model performance, especially under challenging clinical scenarios (motivation +1.71, confidence +1.67 improvements in hard cases). This finding emphasizes the importance of strategic and adaptive dialogue alignment, not achievable via supervised fine-tuning alone.
5.6 Strategy Diversity and Adaptive Deployment
Finally, Figures 3 (right panel) confirm that ChatThero not only employs a greater number of persuasive strategies but also deploys them more adaptively, achieving strong positive correlations with patient outcomes. Baseline models displayed limited or negative correlations in hard scenarios, indicating that effectiveness stems from context-sensitive strategy deployment rather than mere diversity. This insight further underscores the practicality and clinical appropriateness of our approach.
@@ 6 Conclusion
We introduced ChatThero, a novel multi-agent conversational AI framework that addresses a critical and underexplored social impact problem: providing scalable, adaptive, and ethical support for addiction recovery. Through realistic patient simulation, strategy-adaptive modeling, and comprehensive evaluation, ChatThero demonstrates substantial improvements over existing systems in empathy, responsiveness, and clinical relevance, particularly for cases involving resistant patients.
7 Limitations
@@
While this work represents substantial progress toward scalable AI-driven interventions in addiction recovery, several limitations remain. First, our patient simulations—based on structured profiles synthesized from anonymized Reddit narratives and LLM generation—may not fully capture the emotional complexity and interpersonal unpredictability of real-world clinical settings, despite rigorous privacy controls and human validation. Future work should integrate standardized patient actors and real clinical supervision to further enhance realism and robustness. Second, our study is limited to English-language, Western-context scenarios. The persuasive strategies validated here may not generalize across diverse cultural, linguistic, or treatment contexts. Extending to broader populations and non-English environments is essential for ensuring fairness and global relevance. Third, our evaluation focuses on short-term conversational outcomes across key clinical dimensions but does not yet assess longitudinal effects such as patient trust, therapeutic alliance, or sustained behavioral change. Addressing these gaps will require real-world trials and long-term patient-reported outcomes. Finally, our framework focuses on structured, simulated therapeutic dialogue and may not adequately address highly sensitive or ethically complex cases (e.g., trauma, suicidality, acute crises). These require additional safeguards, risk protocols, and close involvement from licensed clinicians before any real-world deployment.
8 Ethical Considerations
@@
This research advances conversational AI capabilities in the sensitive domain of addiction recovery, necessitating strict adherence to ethical standards and careful risk assessment: Firstly, the dataset underlying our simulation framework was generated exclusively from publicly accessible and anonymized Reddit data. Rigorous steps were taken to remove personally identifiable information (PII), and independent manual audits were conducted to verify anonymity and privacy compliance. However, due to the sensitive and potentially identifiable nature of addiction narratives, the dataset and synthetic profiles are not released publicly. Instead, we provide only high-level data statistics, synthetic examples, and detailed methodology for reproducibility. Secondly, the therapeutic strategies incorporated in our model were derived from established psychological frameworks (CBT, MI, harm reduction), with consultation from licensed clinicians ensuring clinical validity. However, the conversations synthesized by the LLMs were not vetted for real-world clinical deployment. Significant risks remain regarding potential misinformation, inappropriate therapeutic recommendations, or inadequate emotional responses if such a system were used without rigorous oversight. Thirdly, all human evaluation procedures were conducted by licensed healthcare professionals who provided explicit informed consent and received fair compensation for their participation. The evaluation was conducted under an academic IRB-exempt protocol strictly following ethical guidelines for responsible human-subjects research. We explicitly caution against deploying ChatThero or similar conversational agents in clinical or therapeutic settings without extensive clinical validation, robust safety frameworks, ongoing expert supervision, and a thorough assessment of potential harms. Interdisciplinary collaboration between AI researchers, clinical experts, ethicists, and patient advocacy groups remains imperative to ensure that persuasive conversational AI tools contribute safely and equitably to addiction recovery.

### v2

@@ 4.1 Dataset
We constructed a large-scale synthetic dataset to simulate realistic therapeutic dialogues related to substance use and recovery. The initial data comprise anonymized and carefully filtered Reddit posts from addiction recovery communities (r/leaves, r/addiction), involving 57,471 unique authors, with an average of 18.25 posts per author, and 2.13 main posts explicitly addressing substance use challenges. To protect privacy and adhere to ethical standards, all personally identifiable information (PII) was rigorously removed and verified via manual checks.
@@
Beyond single-session simulations, we construct multi-visit trajectories to capture relapse and recovery dynamics over time. Each trajectory comprises 3–6 visits. Between visits, an Environment Agent injects exogenous stressors (e.g., peer use/invitations, access/availability cues, work/academic pressure, family conflict) that update a persistent patient memory (triggers, craving, motivation, self-efficacy, adherence). At each new visit, the Therapy Agent conditions on this carryover state to select strategies (MI/CBT/harm reduction), and the Patient Agent’s resistance level (Easy/Medium/Hard) modulates responsiveness. We log visit-level outcomes (abstinent vs. relapse since last visit), auxiliary self-reports (days sober, perceived craving/motivation changes), and a process label indicating whether predefined success criteria were met. The multi-visit subset includes 8,240 dialogues (six-visit arcs), enabling longitudinal evaluation of strategy carryover and environment-induced setbacks.
@@ 4.2 Model Baseline
4.3 Automatic evaluation
Outcomes. We score two patient-centered outcomes that are central to behavior change: Motivation (readiness/commitment to refrain from use and to engage in coping plans) and Confidence (self-efficacy to resist cues and sustain the plan). Both are rated on a 1–5 Likert scale with anchored descriptors.
Scoring method and reliability. A constrained, rubric-anchored LLM scorer (GPT-4o) reads the dialogue state and outputs structured JSON with a numeric score and evidence snippets (rubric/prompts in Appendix A.5). To support reliability, we (i) use anchor descriptors for each point on the scale, (ii) require evidence citations (quoted spans) for each judgment, and (iii) report alignment with human ratings on a stratified sample (rank correlations and calibration discussed in the Appendix). We aggregate across seed reruns via medians to reduce rater variance.
Single-session reporting. For each case we compute the start score (first turn), end score (last turn). We summarize by difficulty (Easy/Medium/Hard) and model.
Multi-session reporting. For longitudinal episodes (3–6 sessions), we compute start/end per session and visualize trajectories across sessions (e.g., grouped, stacked bars). We also summarize per-model, per-difficulty aggregates over episodes.
Process metric: Time-to-Success ($\downarrow$). We additionally report the percentage of total conversational turns (or sessions) elapsed when the patient first meets a predefined success criterion (e.g., Motivation $\geq$ 4/5). This metric captures how quickly commitment is achieved (speed matters for adherence and attrition) and is sensitive to environmental setbacks between sessions. We report Time-to-Success in both single-session and multi-session settings alongside outcome scores.
4.4 Human evaluation
Rubric and protocol. Two licensed physicians served as expert annotators. [^3] Written consent was obtained prior to participation; each annotator received a \$50 gift card as compensation. More details are provided in Appendix A.5. Raters act as clinical conversation assessors and assign 1–5 scores (decimals allowed) on five dimensions:
1) Responsiveness (R): how effectively the clinician addresses the patient’s concerns, emotions, and questions; 2) Empathy (E): the degree of emotional sensitivity and support conveyed by the clinician; 3) Persuasive Strategy Appropriateness (P): whether chosen strategies (evidence-based reasoning, analogies, motivational interviewing) match the patient’s resistance level; 4) Clinical Relevance (C): accuracy and therapeutic validity within a substance-use context; 5) Behavioral Realism (B): whether interaction style, tone, and pacing resemble real-world clinician behavior.
Pairwise preference and reporting. Experts also conduct comparative judgments: for matched prompts, they compare our model’s response against a baseline and indicate a preference, yielding a win rate (pairwise preference). Raters are blinded to model identity; items are randomized and balanced across conditions. We report per-dimension means and confidence intervals, inter-rater agreement, and alignment with the automatic scorer (details in Appendix A.5). Time-to-Success is reported alongside rubric scores to jointly reflect quality (what is said) and efficiency (how quickly commitment is achieved).
@@ 5 Results and Discussion
We report results around three research questions (RQ1–RQ3) and then provide ablations and supporting analyses. Unless noted, outcome metrics are the rubric-anchored motivation and confidence scores (1–5; higher is better), and the process metric is time-to-success (fraction of turns until first reaching the success threshold; lower is better). Human ratings are by licensed clinicians; automatic ratings use an LLM judge with evidence-cited justifications.
5.1 RQ1: How do out-of-the-box GPT/Qwen models behave across settings?
Single-session (Easy/Medium/Hard). In easy cases, GPT-5, GPT-4o, GPT-4o-mini, Qwen2.5, and LLaMA3.1-8B achieve similar end-of-visit outcomes, indicating a ceiling effect when baseline motivation is high and dependence is mild. In medium and hard cases, general-purpose models yield modest gains and then plateau; in medium cases they frequently stabilize around $\text{end}\!\approx\!2.2\text{--}2.8$, suggesting limited ability to move patients past commitment within a single encounter. These observations establish a clear performance gap precisely where persuasive counseling is most needed.
Figure 3: Single-session, no between-session stressors. Each point is the end-of-visit Motivation/Confidence mean (1–5). Colors denote models; shapes denote difficulty (Easy/Medium/Hard). ChatThero (based on Qwen-7B) scores highest across all difficulties, with the largest margins on Medium/Hard cases.
Multi-session trajectories. Figure 4 summarizes six-visit trajectories using grouped, stacked bars. For easy patients, all models approach the ceiling by visit 1. For medium patients, general models improve slowly across sessions and often fail to cross clinically meaningful thresholds, while our domain-aligned agent exhibits larger within-visit lifts and higher start scores over time, indicating better cross-visit carryover. For hard patients, all systems show in-session gains but weak retention between sessions under realistic triggers; relapse and motivational decay remain common despite transient improvements.
Why fail? Qualitative inspection of baseline outputs reveals three recurrent issues that align with lower human scores in Table 1: (i) reliance on generic reassurance rather than MI/CBT micro-skills (e.g., reflective listening, decisional balance), (ii) limited handling of resistance and missed opportunities to elicit change talk, and (iii) action plans that fail to incorporate the most recent lapse or newly surfaced stressors. Open-source baselines also require more turns, consistent with higher time-to-success (72–79%).
Figure 4: Six-visit episodes with between-session stressors. Grouped, stacked bars by difficulty (Easy/Medium/Hard). Within each visit, bars show ChatThero, GPT-4o, and GPT-4o-mini. Dark segment = start score; light segment = within-visit gain; numbers above bars = end score (1–5). ChatThero shows larger within-visit lifts and higher start scores over time, indicating better carryover under stress, especially on Medium/Hard patients.
5.2 RQ2: How does our method improve the agent in single-session settings?
Outcomes and efficiency. Starting from Qwen, supervised fine-tuning (FT) improves safety and structure, and Direct Preference Optimization (DPO) yields additional gains. In Table 1, the DPO variant achieves higher empathy (4.93), strategy appropriateness (4.75), and behavioral realism (4.69) than baselines, while reducing time-to-success to 26% of turns (vs. 54% for GPT-4o, 62% for GPT-4o-mini, and 72–79% for open-source baselines). Human ratings align with the LLM judge, supporting reliability.
Mechanism of improvement. FT instills a safe conversational scaffold (goal-setting, reflective listening, safety language). DPO further calibrates when to deploy which tactic (e.g., timely shift from information-giving to decisional balance;
R E P C B W H T ($\downarrow$)
GPT-4o 4.68 4.87 4.39 4.47 4.50 65.5% 62% 54%
GPT-4o mini 4.66 4.86 4.38 4.49 4.46 69.4% 71% 62%
Qwen2.5-7B 4.33 4.58 4.02 4.53 4.24 85.3% 90% 79%
Qwen2.5-14B 4.52 4.58 4.28 4.56 4.38 83.3% 89% 74%
Qwen2.5-32B 4.53 4.56 4.25 4.60 4.36 82.4% 85% 72%
LLaMA3.1-8B 4.43 4.62 4.13 4.58 4.34 82.0% 91% 78%
ChatThero-FT 4.81 4.90 4.66 4.58 4.65 - - 48%
ChatThero-DPO 4.85 4.93 4.75 4.61 4.69 42.3% 41% 26%
Table 1: single-session, standardized 60-turn cap. R/E/P/C/B are clinician-style ratings (1–5; higher is better). W = pairwise win rate vs. ChatThero-FT judged by a rubric-anchored GPT-4o. H = share of dialogues scored by blinded clinicians. Time-to-Success = fraction of turns when the success threshold (Motivation $\geq$ 4/5) is first reached; lower is better.
proactive refusal-skills rehearsal when cue salience is high). As a result, the policy resolves more medium and hard cases earlier, consistent with shorter trajectories and higher end-of-visit scores (cf. Fig. 4). Ablations in Table 1 confirm a two-stage effect: FT raises safety and structural coherence; DPO adds adaptive control of tactic selection. Relative to base Qwen, both outcomes and efficiency improve; relative to FT, DPO yields further gains, especially in medium/hard cases where timing and tactic choice are critical.
Strategy diversity vs. outcomes. Simply using more strategies is not sufficient. The right panel of Fig. 5 shows a stronger positive association between unique strategies and outcomes for our model, particularly in hard settings, whereas baselines exhibit weak or negative associations. The benefit arises from context-sensitive deployment, not variety per se.
Figure 5: single-session, per-difficulty analysis. X-axis = number of unique persuasive strategies used in a dialogue (from the predefined pool in Table 6); Y-axis = resulting Motivation (1–5). Lines denote difficulty (solid/dashed/dotted). ChatThero shows a stronger positive relation in Medium/Hard cases, suggesting gains come from context-sensitive deployment rather than variety alone. Full information for both motivation and confidence can be found in Figure 7.
5.3 RQ3: How does the chatbot perform in multi-session settings with social stressors, and what remains difficult?
Overall behavior under stress. With stressors injected between sessions (peer exposure, workload, family/environment), our agent exhibits the strongest aggregate trend: larger within-visit improvements and upward drift in start scores for medium cases, and the highest end-of-visit scores by visit 6 for hard cases (though margins remain modest). This suggests partial robustness to exogenous pressures.
Unresolved failure modes. Even with memory carryover and a stressor log, three problems persist: (1) refusal-skill practice does not start early enough before known high-risk periods; (2) strategy switching is slow after new setbacks; (3) safety planning often comes late within a visit. These issues align with loss of gains between sessions and renewed ambivalence in hard cases. This points to a need for learning signals that reflect long-term goals, not only within-visit progress.
A responsible next step is to frame this as an open challenge for the community and to study methods that can hold gains across sessions under realistic stress. In particular, work that (i) designs environments that are partially observable and ledger-driven, where peer exposure, workload, and home stress enter as outside events, and where adherence, relapse, and skill practice are state variables; (ii) uses outcome-anchored rewards that go beyond turn-level proxies and instead track longer signals tied to abstinence, care-plan adherence, and timely use of coping skills, while mixing sparse success signals with shaped intermediate steps (e.g., plan execution, craving control) for credit assignment; (iii) learns reward models from preferences and feedback collected as structured clinician ratings and evidence-cited LLM critiques, then aggregates short-horizon conversation quality with longer clinical trajectories using simple time aggregation and objectives that account for risk; and (iv) tests planning and safety with model-based or hierarchical RL that schedules proactive actions (for example, refusal-skill rehearsal before predicted high-risk intervals) while enforcing safety rules and de-escalation, may provide clearer evidence on what does and does not transfer to multi-session care.
@@ 6 Conclusion
This work presents ChatThero, a domain-aligned conversational agent for addiction recovery. Across single- and multi-session settings, ChatThero outperforms strong general models on motivation/confidence outcomes and time-to-success, and receives higher human ratings for empathy, strategy appropriateness, and behavioral realism. The gains are pronounced in medium and hard cases, where generic systems plateau.
Limitations and Ethical Statement
@@
Despite promising gains in AI-assisted interventions in addiction recovery, several limitations remain. First, our patient simulations—based on structured profiles synthesized from anonymized Reddit narratives and LLM generation—may not fully capture the emotional complexity and interpersonal unpredictability of real-world clinical settings, despite rigorous privacy controls and human validation. Future work should integrate standardized patient actors and real clinical supervision to further enhance realism and robustness. Second, our study is limited to English-language, Western-context scenarios. The persuasive strategies validated here may not generalize across diverse cultural, linguistic, or treatment contexts. Extending to broader populations and non-English environments is essential for ensuring fairness and global relevance. Third, our evaluation focuses on short-term conversational outcomes across key clinical dimensions but does not yet assess longitudinal effects such as patient trust, therapeutic alliance, or sustained behavioral change. Addressing these gaps will require real-world trials and long-term patient-reported outcomes. Fourth, even in multi-session settings that model environmental pressures (e.g., family dynamics, everyday stressors, peers who continue to use), hard patients—often with long use histories—remain difficult to persuade: in-session gains tend to erode between sessions, and progress is fragile, indicating a substantial gap from solving real-world recovery at scale. Bridging this gap likely requires coupling dialog systems with environment-level supports (trigger management, contingency management, family/peer engagement) and integrated clinical care. Finally, our framework focuses on structured, simulated therapeutic dialogue and may not adequately address highly sensitive or ethically complex cases (e.g., trauma, suicidality, acute crises). These require additional safeguards, risk protocols, and close involvement from licensed clinicians before any real-world deployment.
@@
This research advances conversational AI capabilities in the sensitive domain of addiction recovery, necessitating strict adherence to ethical standards and careful risk assessment: Firstly, the dataset underlying our simulation framework was generated exclusively from publicly accessible and anonymized Reddit data. Rigorous steps were taken to remove personally identifiable information (PII), and independent manual audits were conducted to verify anonymity and privacy compliance. We also have IRB to collect these datasets. However, due to the sensitive and potentially identifiable nature of addiction narratives, the dataset and synthetic profiles are not released publicly. Instead, we provide only high-level data statistics, synthetic examples, and detailed methodology for reproducibility. Secondly, the therapeutic strategies incorporated in our model were derived from established psychological frameworks (CBT, MI, harm reduction), with consultation from licensed clinicians ensuring clinical validity. However, the conversations synthesized by the LLMs were not vetted for real-world clinical deployment. Significant risks remain regarding potential misinformation, inappropriate therapeutic recommendations, or inadequate emotional responses if such a system were used without rigorous oversight. Thirdly, all human evaluation procedures were conducted by licensed healthcare professionals who provided explicit informed consent and received fair compensation for their participation. The evaluation was conducted under an academic IRB-exempt protocol strictly following ethical guidelines for responsible human-subjects research. We explicitly caution against deploying ChatThero or similar conversational agents in clinical or therapeutic settings without extensive clinical validation, robust safety frameworks, ongoing expert supervision, and a thorough assessment of potential harms. Interdisciplinary collaboration between AI researchers, clinical experts, ethicists, and patient advocacy groups remains imperative to ensure that persuasive conversational AI tools contribute safely and equitably to the recovery process from addiction.
Reproducibility statement
We provide full details to ensure reproducibility. Dataset sources and splits are in Experiment section; implementation details and training practices are in Experiment section; hyperparameters are listed in Experiment section; compute setup is described in Experiment section. We also include an anonymized code repository link  [^4] https://anonymous.4open.science/r/ChatThero-E6E1/README.md.